OpenStudy (anonymous):

What is the similarity ratio of a prism with surface area 36 ft2 to a similar prism with surface area 225 ft2?

OpenStudy (anonymous):

answer choice?

OpenStudy (jdoe0001):

\(\bf \cfrac{area1}{area2}=\cfrac{(side1)^2}{(side2)^2}\implies \cfrac{area1}{area2}=\left(\cfrac{side1}{side2}\right)^2 \\ \quad \\ \sqrt{\cfrac{area1}{area2}}=\cfrac{side1}{side2}\iff\textit{similarity ratio}\qquad thus \\ \quad \\ \cfrac{36}{225}=\left(\cfrac{\square }{\square }\right)\implies \sqrt{\cfrac{36}{225}}=\cfrac{\square }{\square }\implies \cfrac{\sqrt{36}}{\sqrt{225}}=\cfrac{\square }{\square }\)

OpenStudy (jdoe0001):

hmm \(\bf \cfrac{area1}{area2}=\cfrac{(side1)^2}{(side2)^2}\implies \cfrac{area1}{area2}=\left(\cfrac{side1}{side2}\right)^2 \\ \quad \\ \sqrt{\cfrac{area1}{area2}}=\cfrac{side1}{side2}\iff\textit{similarity ratio}\qquad thus \\ \quad \\ \cfrac{36}{225}=\left(\cfrac{\square }{\square }\right)^2\implies \sqrt{\cfrac{36}{225}}=\cfrac{\square }{\square }\implies \cfrac{\sqrt{36}}{\sqrt{225}}=\cfrac{\square }{\square }\)

OpenStudy (anonymous):

6.25